# Thermodynamics of Schwarzschild-AdS black hole in non-commutative geometry

### v3

@@ I Introduction
Non-commutative geometry is a theory of spacetime quantization noncomu1 ; noncomu2 ; noncomu3 ; noncomu4 ; noncomu5 ; noncomu6 ; noncomu7 , described by commutator $\left[\hat{x}^{\mu},\hat{x}^{\nu}\right]=\mathrm{i}\hat{\Theta}^{\mu\nu}$, where $\hat{x}^{\mu}$ and $\hat{x}^{\nu}$ are spacetime coordinate operators, $\hat{\Theta}^{\mu\nu}$ is an anti-symmetric constant tensor, $\Theta$ is a parameter with dimension of $\left[\mathrm{L}^{2}\right]$ that represents the minimum spacetime scale, and $\sqrt{\Theta}$ is a small quantity of the order of the Planck length $\ell_{p}$ commutator ; planck-length . Extensive research has been conducted to consider the effects of non-commutative fields in gravitational theories gr1 ; gr2 ; gr3 ; gr4 ; gr5 ; gr6 ; gr7 ; gr8 . An interesting fact is that the commutator of coordinate operators being a nonzero constant $\Theta$ matrix is incompatible with Lorentz covariance noncomu1 , which potentially suggests a close relationship between non-commutative fields and the violation of Lorentz symmetry. On the other hand, this also provides a new insight into theories about the violation of Lorentz covariance lorentz1 ; lorentz2 ; lorentz2 ; lorentz3 ; lorentz4 ; lorentz5 ; lorentz6 . Non-commutative geometry also provides a reasonable modification to Schwarzschild spacetime. Compared to the point-like mass distribution in the ordinary Schwarzschild spacetime, the mass distribution of a spherically symmetric star is altered to a Gaussian distribution $\rho=\frac{M}{\left(4\pi\Theta\right)^{\frac{3}{2}}}\exp\left(-\frac{r^{2}}{4\Theta}\right)$ and a Lorentzian distribution $\rho=\frac{M\sqrt{\Theta}}{\pi^{\frac{3}{2}}\left(r^{2}+\pi\Theta\right)^{2}}$ spread throughout the entire space distribution1 . Unlike the Schwarzschild spacetime with singularities, these two distributions both lead to a regular de Sitter core at short distances. There have been many studies investigating the properties of black holes (BHs) in the background of non-commutative geometry distribution1 ; gr1 ; gr2 ; gr3 ; gr4 ; Gaussianthermo ; optics ; thermo1 ; thermo2 ; thermo3 . In gr1 , the authors found that the evaporation endpoint of a Schwarzschild black hole, corrected by a Gaussian distribution, is identified as a zero-temperature extremal black hole, even in the case of electrically neutral and non-rotating objects. Moreover, there exists a finite maximum temperature that the BH can attain before eventually cooling down to absolute zero. The Schwarzschild BH with the Lorentzian distribution has been extensively studied as well, including its gravitational lensing effects, quasinormal modes, BH shadows, and thermodynamic properties optics .
@@ II Schwarzschild-AdS BH in non-commutative geometry
The Lorentzian distribution $\rho$ of mass density in non-commutative geometry is distribution1 ; optics ; gr1 ; Gaussianthermo
@@
$f\left(r\right)=1-\frac{C}{r}+\frac{8M\sqrt{\Theta}}{\sqrt{\pi}r^{2}}-\frac{\Lambda r^{2}}{3}+\mathcal{O}\left(\Theta^{\frac{3}{2}}\right),$ (5)
where $C$ is integration constant. To insure that this metric can degenerate to the Schwarzschild-AdS spacetime when $\Theta$ goes to zero, $C$ should be chosen as $2M$. It is clear that the correction of non-commutative geometry is of higher order.
@@
$M>a,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ r_{h}>\frac{a}{2}.$ (8)
@@ III Thermodynamics
III.1 Thermodynamic functions and the first law
@@
where $A=\!\displaystyle\int\!\!\!\!\int\sqrt{g_{\theta\theta}g_{\phi\phi}}\mathrm{d}\theta\mathrm{d}\phi=4\pi r_{h}^{2}$ is the surface area of the BH’s event horizon. The above functions are helpful for deriving the thermodynamic first law. But, when the energy-momentum tensor $T_{0}^{0}$ includes the BH’s mass (as seen in Eq. 4), the conventional first law of thermodynamics $\mathrm{d}M=T\mathrm{d}S+V\mathrm{d}P$ will be violated thermo_4 ; rev1 ; rev2 . It’s easy to verify
@@
$\mathrm{d}\mathcal{M}=W\mathrm{d}M=T\mathrm{d}S+V\mathrm{d}P,$ (14)
@@
$T=W\left(\frac{\partial M}{\partial S}\right)_{P},$ (16)
@@
$V=W\left(\frac{\partial M}{\partial P}\right)_{S}=\frac{4\pi r_{h}^{3}}{3}.$ (17)
@@
III.2 Critical point
To investigate the criticality, the specific volume $v=2r_{h}$ is introduced RNcriticality ; v . By using this relation, Eq. 9 and Eq. 10 will give the equation of state
@@ III.2 Critical point
$v_{c}=4.8831a,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ T_{c}=\frac{0.036403}{a},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ P_{c}=\frac{0.0027338}{a^{2}}.$ (21)
@@
This value is smaller than that of the Van der Waals system (0.375) RNcriticality . The graph of $P\left(v\right)$ is plotted in Fig. 1.
@@
As seen in the figure, there is only one critical point for a fixed parameter $a$. For an isothermal process $T<T_{c}$, the small BH-large BH phase transition occurs. The part where the pressure increases is unstable, and useful solution is to introduce an isobaric curve, which represents the process of the phase transition. The position of this isobaric curve is determined by Maxwell’s area law RNcriticality ; thermo_4 ; maxareapv1 ; maxareapv2
@@
$p=\frac{P}{P_{c}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tau=\frac{T}{T_{c}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \nu=\frac{v}{v_{c}},$ (24)
@@
$t=\frac{T}{T_{c}}-1=\tau-1,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ w=\frac{V}{V_{c}}-1=\nu^{3}-1.$ (26)
@@
There are four critical exponents exp1 :
$C_{V}=T\left(\frac{\partial S}{\partial T}\right)_{V}\propto\left|t\right|^{-\alpha},$ (28)
$\Delta V\propto\left|t\right|^{\beta},$ (29)
@@
$\left|P-P_{c}\right|\propto\left|V-V_{c}\right|^{\delta},$ (31)
@@
It is concluded that four critical exponents of the Schwarzschild-AdS BH in non-commutative geometry are the same as those of the Van der Waals system RNcriticality . These four exponents conform to the Griffiths, Rushbrooke and Widom functions maxareapv1 ; exp1 ; exp2
$\text{Griffiths}\!:\leavevmode\nobreak\ \alpha+\beta\left(\delta+1\right)-2=0,$ (37)
$\text{Griffiths}\!:\leavevmode\nobreak\ \gamma\left(\delta+1\right)+\left(\alpha-2\right)\left(\delta-1\right)=0,$ (38)
$\text{Rushbrooke}\!:\leavevmode\nobreak\ \alpha+2\beta+\gamma-2=0,$ (39)
$\text{Widom}\!:\leavevmode\nobreak\ \gamma-\beta\left(\delta-1\right)=0.$ (40)
III.3 Isobaric heat capacity
@@
III.4 Gibbs free energy and zeroth-order phase transition
@@
Considering Maxwell’s area law in $\left(S,T\right)$ coordinate system maxareats
@@
The Gibbs free energy, corrected by Maxwell’s area law, is plotted in Fig. 8.
@@
Here, a peculiar phenomenon occurs: according to Ehrenfest’s classification of phase transitions, a discontinuous Gibbs free energy indicates the occurrence of a zeroth-order phase transition. The zeroth-order phase transition of BHs is a very intriguing phenomenon. Studies have shown that the dilaton– electromagnetic coupling interaction could lead to a zeroth-order phase transition maxareapv2 ; zero1 ; zero2 . However, unlike the physical mechanism that induces zeroth-order phase transitions as described in maxareapv2 ; zero1 ; zero2 , our research indicates that for a static, spherically symmetric AdS BH, the violation of the conventional first law of thermodynamics would lead to a change in the Gibbs free energy of a BH after undergoing a small BH-large BH phase transition.
@@ IV Joule-Thomson process
The Joule-Thomson expansion in BH thermodynamics has been widely studied JT1 ; JT2 ; JT3 ; JT4 ; JT5 ; JT6 ; JT7 ; JT8 . The Joule-Thomson process is a thermodynamic system’s throttling (constant enthalpy) expansion. As previously mentioned, the mass $M$ of an AdS-BH is considered as its enthalpy. Even though the first law has been modified, studying the constant mass process of BHs remains highly significant.
@@
This value lies between the results of the RN-AdS BH (0.5) and the Van der Waals system (0.75). And, interestingly, study had shown that the Joule-Thomson expansion of the Van der Waals systems reveals the presence of a maximum inversion temperature $T_{i}^{max}$ JT1 , whereas the BH we are studying does not exhibit this phenomenon.
@@ Acknowledgments
We want to thank the National Natural Science Foundation of China (Grant No.11571342) for supporting us on this work.
@@ Appendix A Differential form $\mathrm{d}\mathcal{M}$ is not an exact form
This form is exact if and only if
@@
$T=W\left(\frac{\partial M}{\partial S}\right)_{P},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ V=W\left(\frac{\partial M}{\partial P}\right)_{S}.$ (A.3)
@@
$\left(\frac{\partial W}{\partial S}\right)_{M}\left(\frac{\partial S}{\partial M}\right)_{P}=0.$ (A.5)
@@
So $\mathrm{d}\mathcal{M}$ has been proven to not be an exact form. Thus, form $-S\mathrm{d}T+V\mathrm{d}P$ is not exact either, because if there exists a function $\mathcal{G}$ satisfying $\mathrm{d}\mathcal{G}=-S\mathrm{d}T+V\mathrm{d}P$, then one can construct $\mathcal{M}=\mathcal{G}+TS$ to meet $\mathrm{d}\mathcal{M}=T\mathrm{d}S+V\mathrm{d}P$, which causes a contradiction.

### v5

@@ I Introduction
Non-commutative geometry is a theory of spacetime quantization [1, 2, 3, 4, 5, 6, 7], described by commutator $\left[\hat{x}^{\mu},\hat{x}^{\nu}\right]=\mathrm{i}\hat{\Theta}^{\mu\nu}$, where $\hat{x}^{\mu}$ and $\hat{x}^{\nu}$ are spacetime coordinate operators, $\hat{\Theta}^{\mu\nu}$ is an anti-symmetric constant tensor, $\Theta$ is a parameter with dimension of $\left[\mathrm{L}^{2}\right]$ that represents the minimum spacetime scale, and $\sqrt{\Theta}$ is a small quantity of the order of the Planck length $\ell_{p}$ [8, 9]. Extensive research has been conducted to consider the effects of non-commutative fields in gravitational theories [10, 11, 12, 13, 14, 15, 16, 17]. An interesting fact is that the commutator of coordinate operators being a nonzero constant $\Theta$ matrix is incompatible with Lorentz covariance [1], which potentially suggests a close relationship between non-commutative fields and the violation of Lorentz symmetry. On the other hand, this also provides a new insight into theories about the violation of Lorentz covariance [18, 19, 19, 20, 21, 22, 23]. Non-commutative geometry also provides a reasonable modification to Schwarzschild spacetime. Compared to the point-like mass distribution in the ordinary Schwarzschild spacetime, the mass distribution of a spherically symmetric star is altered to a Gaussian distribution $\rho=\frac{M}{\left(4\pi\Theta\right)^{\frac{3}{2}}}\exp\left(-\frac{r^{2}}{4\Theta}\right)$ and a Lorentzian distribution $\rho=\frac{M\sqrt{\Theta}}{\pi^{\frac{3}{2}}\left(r^{2}+\pi\Theta\right)^{2}}$ spread throughout the entire space [24]. Unlike the Schwarzschild spacetime with singularities, these two distributions both lead to a regular de Sitter core at short distances. There have been many studies investigating the properties of black holes (BHs) in the background of non-commutative geometry [24, 10, 11, 12, 13, 25, 26, 27, 28, 29]. In [10], the authors found that the evaporation endpoint of a Schwarzschild black hole, corrected by a Gaussian distribution, is identified as a zero-temperature extremal black hole, even in the case of electrically neutral and non-rotating objects. Moreover, there exists a finite maximum temperature that the BH can attain before eventually cooling down to absolute zero. The Schwarzschild BH with the Lorentzian distribution has been extensively studied as well, including its gravitational lensing effects, quasinormal modes, BH shadows, and thermodynamic properties [26].
@@ II Schwarzschild-AdS BH in non-commutative geometry
The Lorentzian distribution $\rho$ of mass density in non-commutative geometry is [24, 26, 10, 25]
@@
$f\left(r\right)=1-\frac{2M}{r}+\frac{8M\sqrt{\Theta}}{\sqrt{\pi}r^{2}}-\frac{\Lambda r^{2}}{3}+\mathcal{O}\left(\Theta^{\frac{3}{2}}\right).$ (5)
It is clear that the correction of non-commutative geometry is of higher order.
@@
$M>a,~{}~{}~{}r_{h}>\frac{a}{2}.$ (8)
@@ III Thermodynamics
The first law and Smarr relation
$M=\frac{-3r_{h}^{2}+\Lambda r_{h}^{4}}{3a-6r_{h}}.$ (9)
@@
where $A=\!\displaystyle\int\!\!\!\!\int\sqrt{g_{\theta\theta}g_{\phi\phi}}\mathrm{d}\theta\mathrm{d}\phi=4\pi r_{h}^{2}$ is the surface area of the BH’s event horizon. The above functions are helpful for deriving the thermodynamic first law. But, when the energy-momentum tensor $T_{0}^{0}$ includes the BH’s mass (as seen in Eq. 4), the conventional first law of thermodynamics $\mathrm{d}M=T\mathrm{d}S+V\mathrm{d}P$ will be violated [37, 38, 41]. It’s easy to verify
@@
$\mathrm{d}\mathcal{M}=W\mathrm{d}M=T\mathrm{d}S+V\mathrm{d}P+\Phi_{a}\mathrm{d}a,$ (14)
@@
$T=W\left(\frac{\partial M}{\partial S}\right)_{P,a},$ (16)
@@
$V=W\left(\frac{\partial M}{\partial P}\right)_{S,a}=\frac{4\pi r_{h}^{3}}{3}.$ (17)
And
$\Phi_{a}=W\left(\frac{\partial M}{\partial a}\right)_{S,P}=\frac{-3r_{h}+\Lambda r_{h}^{3}}{6a-12r_{h}}$ (18)
could be interpreted as ”non-commutative potential”. Treating the noncommutative parameter as a thermodynamic variable is not merely a formal requirement but also facilitates the derivation of the black hole’s Smarr relation.
For BH’s mass function $M=M\left(S,P,a\right)$, performing a scale transformation $r_{h}\rightarrow kr_{h}$ on the system, one could get
$k^{d_{M}}M=M\left(k^{d_{S}}S,k^{d_{P}}P,k^{d_{a}}a\right),$ (19)
where $d_{M}$, $d_{S}$, $d_{P}$ and $d_{a}$ are scale dimensions of $M$, $S$, $P$ and $a$. Taking the derivative of both sides of the above equation with respect to $k$, one obtain
$d_{M}k^{d_{M}-1}M=M_{S}d_{S}k^{d_{S}-1}S+M_{P}d_{P}k^{d_{P}-1}P+M_{a}d_{a}k^{d_{a}-1}a,$ (20)
where $M_{S}=\frac{\partial M}{\partial S}$. The same applies to $P$ and $a$. One could set $k=1$ and derive
$d_{M}M=M_{S}d_{S}S+M_{P}d_{P}P+M_{a}d_{a}a.$ (21)
According to Eqs. 5, 11, 12, $d_{M}=1,d_{S}=2,d_{P}=-2$, $d_{a}=1$. So one finally obtain
$M=2\left(M_{S}S-M_{P}V\right)+M_{a}a.$ (22)
With the help of Eqs. 16, 17, 18, one have
$WM=2\left(TS-PV\right)+\Phi_{a}a.$ (23)
It is clear that when $W=1$, the conventional Smarr relation appears. This relation is general. One only needs to fully consider quantities with scale dimensions and derive the Smarr relation using the same method. The general Smarr relation is follows
$WM=2\left(TS-PV\right)+d_{y^{i}}\Phi_{y^{i}}y^{i},$ (24)
Here, $y^{i}$ represents the $i$-th thermodynamic variable other than entropy and pressure, $d_{y^{i}}$ denotes its scale dimension, $\Phi_{y^{i}}=W\frac{\partial M}{\partial y^{i}}$, and repeated indices imply summation.
@@
Critical point
To investigate the criticality, the specific volume $v=2r_{h}$ is introduced [32, 42]. By using this relation, Eq. 10 and Eq. 11 will give the equation of state
@@
$v_{c}=4.8831a,~{}~{}~{}T_{c}=\frac{0.036403}{a},~{}~{}~{}P_{c}=\frac{0.0027338}{a^{2}}.$ (28)
@@
This value is smaller than that of the Van der Waals system (0.375) [32]. The graph of $P\left(v\right)$ is plotted in Fig. 1.
@@
As seen in the figure, there is only one critical point for a fixed parameter $a$. For an isothermal process $T<T_{c}$, the small BH-large BH phase transition occurs. The part where the pressure increases is unstable, and useful solution is to introduce an isobaric curve, which represents the process of the phase transition. The position of this isobaric curve is determined by Maxwell’s area law [32, 37, 43, 44]
@@
$p=\frac{P}{P_{c}},~{}~{}~{}\tau=\frac{T}{T_{c}},~{}~{}~{}\nu=\frac{v}{v_{c}},$ (31)
@@
$t=\frac{T}{T_{c}}-1=\tau-1,~{}~{}~{}w=\frac{V}{V_{c}}-1=\nu^{3}-1.$ (33)
@@
There are four critical exponents [45]:
$C_{V}=T\left(\frac{\partial S}{\partial T}\right)_{V}\propto\left|t\right|^{-\alpha},~{}$ (35)
$\Delta V\propto\left|t\right|^{\beta},~{}\qquad\quad$ (36)
@@
$\left|P-P_{c}\right|\propto\left|V-V_{c}\right|^{\delta},~{}\quad$ (38)
@@
It is concluded that four critical exponents of the Schwarzschild-AdS BH in non-commutative geometry are the same as those of the Van der Waals system [32]. These four exponents conform to the Griffiths, Rushbrooke and Widom functions [43, 45, 46]
$\text{Griffiths}\!:~{}\alpha+\beta\left(\delta+1\right)-2=0,$ (44)
$\text{Griffiths}\!:~{}\gamma\left(\delta+1\right)+\left(\alpha-2\right)\left(\delta-1\right)=0,$ (45)
$\text{Rushbrooke}\!:~{}\alpha+2\beta+\gamma-2=0,$ (46)
$\text{Widom}\!:~{}\gamma-\beta\left(\delta-1\right)=0.$ (47)
Isobaric heat capacity
@@
Gibbs free energy and zeroth-order phase transition
@@
Considering Maxwell’s area law in $\left(S,T\right)$ coordinate system [40]
@@
When considering Maxwell’s area law, the Gibbs free energy is shown by Fig. 8.
Figure 8: The graph of the Gibbs free energy $G\left(T\right)$ when considering Maxwell’s area law. The points $\left(T_{s},G_{s}\right)$ and $\left(T_{l},G_{l}\right)$ represent the small BH and the large BH. We set $a=1$ and $P=0.4P_{c}$.
In this case, the Gibbs free energy for different pressures is plotted in Fig. 9.
@@
Here, a peculiar phenomenon occurs: according to Ehrenfest’s classification of phase transitions, a discontinuous Gibbs free energy indicates the occurrence of a zeroth-order phase transition. The zeroth-order phase transition of BHs is a very intriguing phenomenon. Studies have shown that the dilaton– electromagnetic coupling interaction could lead to a zeroth-order phase transition [44, 47, 48]. However, unlike the physical mechanism that induces zeroth-order phase transitions as described in [44, 47, 48], our research indicates that for a static, spherically symmetric AdS BH, the violation of the conventional first law of thermodynamics would lead to a change in the Gibbs free energy of a BH after undergoing a small BH-large BH phase transition.
@@ IV Joule-Thomson process
The Joule-Thomson expansion in BH thermodynamics has been widely studied [33, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. The Joule-Thomson process is a thermodynamic system’s throttling (constant enthalpy) expansion. As previously mentioned, the mass $M$ of an AdS-BH is considered as its enthalpy. Even though the first law has been modified, studying the constant mass process of BHs remains highly significant.
@@
This value lies between the results of the RN-AdS BH (0.5) and the Van der Waals system (0.75). And, interestingly, study had shown that the Joule-Thomson expansion of the Van der Waals systems reveals the presence of a maximum inversion temperature $T_{i}^{max}$ [33], whereas the BH we are studying does not exhibit this phenomenon.
@@ Acknowledgments
We want to thank Yu-Cheng Tang and Yu-Hang Feng for their useful suggestions.
Data availability
This paper is a theoretical study. No data was used or generated during the research.
@@ Appendix A Differential form $\mathrm{d}\mathcal{M}$ is not an exact form
Here, we have not explicitly written out the differential terms of the non-commutative parameter $\mathrm{d}a$, as this does not affect the proof. This form is exact if and only if
@@
$T=W\left(\frac{\partial M}{\partial S}\right)_{P},~{}~{}~{}V=W\left(\frac{\partial M}{\partial P}\right)_{S}.$ (A.3)
@@
$\left(\frac{\partial W}{\partial S}\right)_{M}\left(\frac{\partial M}{\partial P}\right)_{S}=0.$ (A.5)
@@
So $\mathrm{d}\mathcal{M}$ has been proven to not be an exact form. There is another method to prove this fact. If there is a scalar function $\mathcal{M}$ satisfying Eq. 14, one could obtain its Smarr relation
$\mathcal{M}=2\left(TS-PV\right)+\Phi_{a}a.$ (A.9)
Then one could get
$\mathcal{M}=WM,$ (A.10)
which leads to
$\mathrm{d}\mathcal{M}=M\mathrm{d}W+W\mathrm{d}M.$ (A.11)
On the other hand, we require that $\mathrm{d}\mathcal{M}=W\mathrm{d}M$. It gives
$\mathrm{d}W=0.$ (A.12)
This doesn’t hold unless $W$ is a constant.
Form $-S\mathrm{d}T+V\mathrm{d}P$ is not exact either, because if there exists a function $\mathcal{G}$ satisfying $\mathrm{d}\mathcal{G}=-S\mathrm{d}T+V\mathrm{d}P$, then one can construct $\mathcal{M}=\mathcal{G}+TS$ to meet $\mathrm{d}\mathcal{M}=T\mathrm{d}S+V\mathrm{d}P$, which causes a contradiction.